3.)calculate the molality of phosphoric acid, H3PO4, in a solution of 29.0g H3PO4 in 250.0g H2O. (atomic mass :H=1.008g/mol:P=30.97g/mol:O=16.00g/mol:Ans.m=1.18mol solute /kg solvent)



4.)Refer to the data given in problem #3, calculate the mole fraction of H3PO4 and H20. (Ans. X H3PO4 =0.979:XH20=0.0209



5.)An aqueous solution has 0.0400g of NaCl (salute) in 2.00kg H2O(solvent). What is the concentration of the solution in parts per million



6.)calculate the volume, in mL, of 1.72M HCl solution that will react with 2.67 mol of CaCO3? (Ans. 3.10x103mL)


__HCl+__CaCl2+__H2O+__CO2



7.)If 0.650% solution of Na2CO3 will be used to precipitate Ca+2 from solution, determine the amount of solution, in grams, needed to precipitate 5.00x102mL of 0.01120M Ca+? (Atomic mass :Na 22.99g/mol:C=12.01g/mol:O=16.00g/mol:Ca=40.08g/mol:Ans.91.3g



___Na2CO3(aq)+__Ca+2(aq)___CaCO3(s)+___Na+(aq)

1
Expert's answer
2022-05-30T17:02:08-0400

3.)

Concentration of molality = solute(mol) / solvent(kg)

n (H3PO4) = m / Mr = 29 / 98 = 0,296 mol

n — amount of substance (mol);

m — mass of substance (g);

Mr — relative molecular mass (g/mol)

Molality= 0,296 mol / 0,25 kg =1,184 mol/kg

ANSWER: 1,184 mol/kg


4.)

n (H3PO4) = 0,296 mol

n (H2O) = m / Mr = 250 / 18 = 13,89 mol

n (total) = 0,296 + 13,89 = 14,185 mol

X (H3PO4) =0,296 mol / 14,185 mol = 0,0209

X (H2O) = 13,89 mol / 14,185 mol = 0,9791

ANSWER:

H2O : 0,09791

H3PO4 : 0,0209


5.)

the density (p) of water is 1 g/ml.

m (NaCl) = 0,04 g = 40 mg

V (H2O) = m * p = 2kg * 1 = 2 litr

ppm = m (mg) / V (l) = 40 / 2 = 20


ANSWER: ppm= 20


6).

CaCO3 + 2 HCl = CaCl2 + CO2 + H2O

According to the above reaction,

1 mol CaCO3 —> 2 mol HCl

2,76mol CaCO3 —> X

X = 2,76 * 2 / 1 = 5,52 mol HCl


V = n / CM = 5,52mol / 1,72mol/l = 3,21 litr or 3,21 * 10ml

V — volume of solution (litr);

n — amount of substance (mol);

CM — concentration of molarity.

ANSWER: 3,21 * 103 ml.


7).

a)

V — volume = 5*102ml = 0,5 litr;

CM — concentration of molarity =

= 0,0112 mol/l;

n — amount of substance (mol) = ?


n (Ca2+) = V * CM = 0,5 * 0,0112 =0,0056 mol

b)

Mr (Na2CO3) = 23*2+12+16*3 = 106 g/mol

Na2CO3 + Ca2+ => CaCO3 + 2Na+

According to the above reaction,

106 g Na2CO3 — > 1mol Ca2+

X — > 0,0056mol Ca2+

X = 106 * 0,0056 / 1 = 0,5936 g Na2CO3

c)

0,5936 g — 0,65%

X — 100%

X = 0,5936 * 100 / 0,65 = 91,3 g

ANSWER: 91,3 g